RC meeting and Run Plan, Tuesday Sept. 26, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time: Monday - Friday at 3:30pm

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


Hall A call-in instructions:

  • Call RC first.
  • Techs on site 7-4:30 Mon-Th, 7-11AM Fr. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
  • After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

Hall Access procedure:

  • Access check list (binder on the counting room desk) needs to be filled out before making an access.
  • Scan (take pictures with phone) and post it to the logbook.
  • Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

New instruction for shift crew when changing target:

  • Beam restoration instructions have been updated. Printed version is in the counting room.
  • When we change target, make sure the SOFs of Target BPMs are set to the correct values.
  • SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for "No Target" has been updated and they are same as the He3 target values)

Status Update

Overall Summary

After fixing the bigbite power supply yesterday https://logbooks.jlab.org/entry/4187236, bigbite didn't trip again during the swing and owl shift.

We went to beam permit just before RC meeting yesterday, and followed the beam recovery procedure. Shift leader had issue with DAQ while taking carbon hole run, and GEM expert helped to fix it remotely. Beam checkout is completed around 6pm and we started production run on He3 @42.2% polarization. We continued production run until around 3:20 am. After that MCC stopped sending beam for BSY corrector testing, and LINAC maintenance.

RadCon started surveying the hall around 9 am this morning, and we are in controlled access now. Subsystem (detector) experts accessed hall to fix the issues related with detectors. Target group accessed the target enclosure for field measurement starting from 1:30 pm (Radcon and hall techs supporting the work).

Accelerator status

  • Maintenance day today
  • Beam recovery during Swing shift

Data collected :

  • He3 data: 0.399 C (corrected charge) [Mon. Swing+ Tue. Owl]
  • accumulated corrected charge on He3 till now: 62.217 C

Hall access

  • SBS GEMs made DAQ unstable (GEM group)
  • Cooling jet hardware FSD limit (Ellen)
  • ECal (Jimmy/Mark Mc.)
  • BBCAL (Provakar)
  • Target group (Gordon/Bill)

Shift schedule

  • Still need to fill open slots


Short Term Run plan

  • Beam recovery after accelerator maintenance
  • Production run on polarized He3 target @35uA beam current (can take higher current?)

==> List of items in case He3 production running isn’t ready:

  1. Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
  2. BigBite Sieve slit => only during day (Techs needed)
    1. with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
    2. 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
  3. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)